OpenStudy (anonymous):

Solve the equation below for x. Round your answer to the nearest thousandth: 10=3^(x+5) -2 A. 5.442 B. -4.558 C. -2.738 D. 2.262

OpenStudy (anonymous):

\[3^{x+5}=12\] \[x+5=\frac{\ln(12)}{\ln(3)}\] \[x=\frac{\ln(12)}{\ln(5)}-5\]

OpenStudy (anonymous):

typo on last one \[x=\frac{\ln(12)}{\ln(3)}-5\]

OpenStudy (anonymous):